y2=7-x2 is this an function
Answer by RAY100(1637)

y^2 = 7-x^2
.
x^2 +y^2 =7
.
This is a circle, center at origin,,(0,0), and radius sqrt 7
.
To be a function, there can be only one value of y for each x value, not so in this case.
.
The other test for a function is the vertical line test,,,when moving a vertical line along the x axis, it should intersect only one value of y,,,,again not so in this case
.
Therefore, we have a relationship, not a function.